OpenStudy (anonymous):

how do i reduce (-√3/2)^2...does the negative remain?

OpenStudy (anonymous):

nope, since it's inside the brackets and the whole thing is squared then the negative disappears and the square root disappears too ^_^ so it's 3/2 again lol

OpenStudy (anonymous):

i thought there was a rule that said if you square a fraction with a square root in the numerator, then the the denominator gets squared too. so in the above example the √3 becomes 3 and the 2 in the bottom become 2^2 = 4...little confused about fractions and exponents

OpenStudy (anonymous):

wait is it sqrt(3) / 2? or sqrt(3/2)?

OpenStudy (anonymous):

it comes from the "unit circle" so its sqrt(3)/2

OpenStudy (anonymous):

oh, then when you square it , it will become 3/4, I thought it was sqrt(3/2) lol sorry :)

OpenStudy (anonymous):

k...perfect...i thought i had it. maybe you know this one though...same question but not sqrt...this one is exponent 100

OpenStudy (anonymous):

e^100?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I used the calculator and got 2.69 x10^43 ^_^

OpenStudy (anonymous):

i know ...crazy value eh...that cant be right

OpenStudy (anonymous):

LOL, why not? as long as it's not negative, then it's right :)

OpenStudy (anonymous):

well this is for using de moivre's theorem to find a complex number....i've yet to see anything that large..then i would multiply that by (cos50pi + i sin50pi).....if you're familiar with this math

OpenStudy (anonymous):

not really, I'm not :( sorry

OpenStudy (anonymous):

that;s cool...thanks for the fraction help...i've been confused about that one all day

OpenStudy (anonymous):

Per Mathematica: \[\left(-\frac{\sqrt{3}}{2}\right)^2=(-1)^2 \left(3^{2/2} 2^{-2}\right)=\frac{3}{4} \]

OpenStudy (anonymous):

you're welcome ^_^